OpenStudy (anonymous):

find two arithmetic means between 64 and 118 (with work if its not too much)

OpenStudy (anonymous):

is this a series?

OpenStudy (anonymous):

i think it is

OpenStudy (whpalmer4):

You're looking for two numbers, \(x,y\), such that \[\frac{x}{64} = \frac{y}{x} = \frac{118}{y} = a\] where \(a\) is the common ratio of the geometric series. That implies that \[x=64*a\]\[y=a*x\]\[118=a*y\]Telescoping them all together\[118=64*a*a*a\]Can you determine the value of \(a\), and from that, the values of the two arithmetic means?

OpenStudy (anonymous):

thanks i actually figured out the answer but that explanation made me more sure! @whpalmer4

OpenStudy (whpalmer4):

What did you get for your answers?

OpenStudy (anonymous):

i got 82 and 100 and the common difference is 18! @whpalmer4

OpenStudy (whpalmer4):

Oh, sorry, you did arithmetic means, not geometric...

OpenStudy (anonymous):

oh no worries!

OpenStudy (whpalmer4):

But, duh, that's what is asked for...I could have sworn I saw geometric!!!

OpenStudy (whpalmer4):

Well, if you need to find geometric means, now you know how :-)